# Personhood credentials: Artificial intelligence and the value of

### v1

@@ B.1   CAPTCHAs and other behavioral filters
Even before AI could directly solve CAPTCHAs, malicious actors found ways to use AI systems to bypass these barriers. For example, testers discovered that GPT-4—a text-only AI model lacking “vision” capabilities—pretended to have a disability to persuade a skeptical human to complete a CAPTCHA on its behalf [198, 176]. More broadly, it is easy for automated scripts to employ workers to solve these tests for fractions of a cent per CAPTCHA.128128128A bot calls on these human workers when it detects it has been presented with a CAPTCHA challenge, through services like 2captcha [70]. For detailed reports, see [271] and [93].
Now, AI systems have developed capabilities that challenge the effectiveness of CAPTCHAs in filtering out such systems. Multimodal AI, which combines vision with cognitive abilities to perform various tasks, in particular, poses a significant challenge.129129129For a review of OpenAI’s GPT-4V and its current abilities to pass certain forms of CAPTCHA, see [195].
@@
It is worth noting that for many service providers that use CAPTCHAs, the aim is not to make bot attacks impossible but to render them economically unviable. CAPTCHA deployers might consider their measures successful if they make abuse more challenging for bots, even if some bots occasionally succeed. Nevertheless, as AI becomes less expensive, the economic barriers imposed by CAPTCHAs diminish, while humans face increased difficulty in solving the more challenging tests.
Some digital services have moved beyond CAPTCHAs to employ a broader range of behavioral filters aimed at reducing bot-based deception. For instance, anomaly detection methods attempt to identify account activity patterns that appear suspiciously coordinated or unusual compared to that of the general user base.130130130For one example of anomalous behavior—unusual signup patterns—from fraudulent users, see [281]. One way that such accounts get discovered is through common patterns in their activity, which may be downstream of a malicious actor needing to program a specific script that executes account signup in automated ways. AI agents might be able to vary their signup practices in ways that are harder to correlate, as opposed to following a relatively static script. JavaScript-based browser challenges assess factors like the manner in which users access a service.131131131For an overview of such challenges and commercial providers of these, see [9]. One way that bots already attempt to appear more person-like is through taking actions like mouse movements and page-scrolling, as well as through time delays. We expect these actions will become more natural-seeming as AI systems learn to better imitate human website behavior, through methods such as behavior cloning [252] or other advanced techniques [121], which can allow for learning complex behavior even when the behavior’s objective cannot be well-specified in advance. For an example of imitating relatively complex human behavior in digital environments, see [18].
These advanced methods fundamentally rely on discernible differences between how people and AI systems interact with a particular website. As AI systems become harder to distinguish from people, websites may face difficult trade-offs. To significantly reduce AI-powered abuse, they might need to limit a substantial portion of people’s activity—which cannot be reliably differentiated from AI-based traffic—in the process. Personhood credentials could offer an alternative that avoids this trade-off by offering a new way to identify automated activity that does not depend on outdated assumptions about AI systems’ cost or sophistication.
@@ B.2   Synthetic content transparency tools
  • Metadata (Provenance):133133133The Partnership on AI uses the term metadata to refer to this mitigation, rather than the term provenance, which sometimes is taken to mean a wider umbrella of mitigations for telling “where content is from” [200]. Metadata, in contrast, refers specifically to “labeling in metadata where the content is from.”including information about content’s origin or editing history, potentially with cryptographic signatures for tamper-resistance (as in C2PA—the most prevalent example of this method.134134134The most prominent metadata provenance group is the Coalition for Content Provenance and Authenticity (C2PA), which aims to “[address] the prevalence of misleading information online through the development of technical standards for certifying the source and history (or provenance) of media content” [61]. C2PA is supported by a number of technology and media companies, including Adobe, Sony, BBC, Google, Microsoft, Meta, and OpenAI, and several technology companies have begun to use its standard for tracking AI-generated images [60, 71]. Similarly, makers of physical cameras, like Sony, and of software for smartphone cameras, like Truepic, are developing tools that enable C2PA (or C2PA-like) tracking so that a photo taken with their equipment would automatically be associated with a C2PA manifest [80, 253].)
  • Fingerprinting: logging AI-generated content into a database so that future content can be checked against this, generally through hashing that compresses the content.135135135This approach has some similarities to hashing illegal sexual material—such as CSAM—to detect it and thus reduce its spread [249].
@@
There are two central reasons why adding personhood credentials can strengthen the overall toolkit of countermeasures for scalable AI-powered deception through synthetic content online.
@@
In some cases, AI-powered abuse can occur without AI-generated content. This could occur when AI leverages human-generated content (Case 1 and Case 2), or when AI interacts directly with websites (Case 3). In other cases, AI-generated content is central to the abuse, but transparency-centric approaches do not fully resolve the harms (Case 4). In yet other cases, AI-generated content is beneficial, and synthetic content transparency tools may be misinterpreted and make these uses less possible, if not complemented by other solutions (Case 5).
@@
Alice wants to find and exploit vulnerabilities on a website to demand an anonymous ransom payment: She uses a wide range of AI agents to analyze the site’s source code and to test out different sequences of unusual inputs. Because the agents interact primarily with the website through mouse-clicks and simple text-field inputs, synthetic content transparency tools are largely unhelpful in this case.142142142Synthetic content transparency tools will not address mouse-clicks and other forms of browser navigation, which are action-based rather than content-based. A website could attempt to look for watermarks in content submitted to its text fields, but watermarking tends to be unreliable for short pieces of text. Moreover, the site receives only a very small portion of content that the AI has generated; it cannot analyze the AI’s private chain-of-thought in planning and executing this abuse.
@@
Bob lost the ability to speak with much fluency after a medical incident, but he was able to recover his ability to speak by using an AI tool. Now, Bob can again speak with others over the phone—even with his own voice. Unfortunately, many companies have started to automatically flag and filter out callers they suspect are voiced by AI, given a surge of AI-powered phone scammers. Because Bob’s voice is in fact AI-generated, he is frequently flagged in these settings and has a harder time getting helpful customer service. Without some alternative method, Bob struggles to prove that he is a real person—using AI to help him express himself—rather than a bad actor using AI to deceive.
@@ C.1   Methods to achieve a one-per-person credential limit
How can an issuer authenticate the existing form of ID, so that they do not accidentally issue personhood credentials based on a spoofed document? Some forms of existing ID are already digitally authenticatable today, whereas others may need to be physically presented in person to avoid spoofing.144144144Some national IDs hold near-field communication (NFC) readable, cryptographically signed information like name, address, and date of birth as well as face images. For a description of NFC in ePassports, see [128]. This allows a holder to demonstrably confirm their ownership of the physical card in the digital domain, unlike ID cards in countries like the United States (including some REAL ID-compliant documents), which often rely upon sending a (spoofable) photo or video for digital authentication (depending on the permissions of the verifying party). One challenge may be how to authenticate IDs from different jurisdictions, in a relatively uniform manner. Usefully, there are already global standards for some forms of ID (such as the United Nations’ oversight of ePassport standards), with public key infrastructure built for authenticating these; third parties can then explore layering zero-knowledge technology atop the documents.145145145For more on ePassports, see [128]. For an overview of how various systems come together to enable identification for cross-border movement, see [73]. For an example of layering zero-knowledge technologies atop these documents, see [287].
@@
Web-of-Trust (WoT) is yet another method for limiting the number of credentials per person: issuers depend on social graph analysis to establish that a person is real (e.g., “Are they vouched for by a person who is strongly believed to be real?”) and has not already received a credential. Often, these systems are “seeded” with initial trusted users from some other method of trust (e.g., existing IDs, biometrics, or in-person “pseudonym parties”)[99]. Someone without a seed-authenticated profile can become authenticated through sufficient vouches from already-trusted parties. While WoT can be a strong method for validating that a person exists, these systems can struggle to confirm uniqueness: A person may be able to get multiple credentials from multiple distinct social circles.152152152Determining whether a person has already received a credential can be a more difficult challenge for WoT-based methods. One way to assess this may be to analyze “Does their social graph look similar to cases where people have been found to have improperly obtained multiple credentials?”, though this will invariably be prone to errors. For a review of WoT-based approaches, see [234]. The most fraud-resistant mechanisms for WoT may involve layering another system, such as biometrics, atop WoT, though in a way such that users still feel comfortable with privacy assurances.153153153For instance, one option for layering simple biometrics atop WoT may include people having a private profile picture, which is viewable only to a small subset of trusted contacts but can still be compared against other private profile pictures in a privacy-preserving way. Use of any biometrics may undermine a system’s reasons for having initially selected WoT as a method, however. To aid in enforcing one-per-person credential limits, WoT issuers may also consider other signals beyond one’s social graph—such as geographic consistency, transaction patterns, and device usage (all commonly used in modern anti-fraud frameworks today).
@@ D.2   Choices for service providers and issuers in an ecosystem with multiple issuers
For a service provider, decisions will often center around which issuers’ PHCs to accept and how these PHC options relate to one another. Trade-offs will often entail balancing the service’s ability to reduce scaled deception by ensuring each person has only one credential, compared with the implications of having fewer choices for its users (such as users feeling pressure to use a credential that is difficult for them to obtain or merely concentrating the power of the chosen issuer, which creates challenges, as discussed in Section 4).

### v2

@@ B.1   CAPTCHAs and other behavioral filters
Even before AI could directly solve CAPTCHAs, malicious actors found ways to use AI systems to bypass these barriers. For example, testers discovered that GPT-4—a text-only AI model lacking “vision” capabilities—pretended to have a disability to persuade a skeptical human to complete a CAPTCHA on its behalf [198, 176]. More broadly, it is easy for automated scripts to enlist human workers to solve these tests for fractions of a cent per CAPTCHA.128128128A bot calls on these human workers when it detects it has been presented with a CAPTCHA challenge, through services like 2captcha [70]. For detailed reports, see [271] and [93].
Now, AI systems have developed capabilities that further challenge the effectiveness of CAPTCHAs in filtering out bots online. Multimodal AI, which combines vision with cognitive abilities to perform various tasks, in particular, poses a significant challenge.129129129For a review of OpenAI’s GPT-4V and its current abilities to pass certain forms of CAPTCHA, see [195].
@@
It is worth noting that for many service providers that use CAPTCHAs, the aim is not to make bot attacks impossible but to render them economically unviable. CAPTCHA deployers might consider their measures successful if they make abuse more challenging for bots, even if some bots occasionally succeed. Nevertheless, as AI becomes less expensive, the economic barriers imposed by CAPTCHAs diminish, while humans might face increased difficulty in solving more challenging tests.
Some digital services have moved beyond CAPTCHAs to employ a broader range of behavioral filters aimed at reducing bot-based deception. For instance, anomaly detection methods attempt to identify account activity patterns that appear suspiciously coordinated or unusual compared to that of the general user base.130130130For one example of anomalous behavior—unusual signup patterns—from fraudulent users, see [281]. One way that such accounts get discovered is through common patterns in their activity, which may be the result of idiosyncratic details in a malicious actor’s automated script. AI agents might be able to vary their practices in ways that are harder to correlate, as opposed to following a relatively static script. JavaScript-based browser challenges assess factors like the manner in which users access a service.131131131For an overview of such challenges and commercial providers of these, see [9]. One way that bots already attempt to appear more person-like is through taking actions like moving the mouse and page-scrolling, as well as through acting on a time delay. We expect these actions will become more natural-seeming as AI systems learn to better imitate human website behavior, through methods such as behavior cloning [252] or other advanced techniques [121], which can allow for learning complex behavior even when the behavior’s objective cannot be well-specified in advance. For an example of imitating relatively complex human behavior in digital environments, see [18].
These advanced methods fundamentally rely on discernible differences between how people and AI systems interact with a particular website. As AI systems become harder to distinguish from people, websites may face difficult trade-offs. To significantly reduce AI-powered abuse, they might need to limit a substantial portion of people’s activity—which cannot be reliably differentiated from AI-based traffic—in the process. Personhood credentials could offer an alternative that avoids this trade-off by offering a new way to identify automated activity that does not depend on outdated assumptions about AI systems’ sophistication.
@@ B.2   Synthetic content transparency tools
  • Metadata (Provenance):133133133The Partnership on AI uses the term metadata to refer to this mitigation, rather than the term provenance, which sometimes is taken to mean a wider umbrella of mitigations for telling “where content is from” [200]. Metadata, in contrast, refers specifically to “labeling in metadata where the content is from.”including information about content’s origin or its edit history, potentially with cryptographic signatures for tamper-resistance (as in C2PA—the most prevalent example of this method134134134The most prominent metadata provenance group is the Coalition for Content Provenance and Authenticity (C2PA), which aims to “[address] the prevalence of misleading information online through the development of technical standards for certifying the source and history (or provenance) of media content” [61]. C2PA is supported by a number of technology and media companies, including Adobe, the BBC, Google, Microsoft, and OpenAI, and several technology companies have begun to use its standard for tracking AI-generated images [60, 71]. Similarly, makers of physical cameras, like Sony, and of software for smartphone cameras, like Truepic, are developing tools that enable C2PA (or C2PA-like) tracking so that a photo taken with their equipment would automatically be associated with a C2PA manifest [80, 253].).
  • Fingerprinting: logging AI-generated content into a database so that future content can be checked against this, generally through hashing.135135135This approach has some similarities to hashing illegal sexual material—such as CSAM—to detect it and thus reduce its spread [249].
@@
There are two central reasons why personhood credentials are an important complementary tool for addressing scalable AI-powered deception online.
@@
In some cases, AI-powered abuse can occur without relying on AI-generated content. This could occur when AI leverages human-generated content (Case 1 and Case 2), or when AI interacts directly with websites (Case 3). In other cases, AI-generated content is central to the abuse, but transparency-centric approaches do not fully resolve the harms (Case 4). In yet other cases, AI-generated content is beneficial, and synthetic content transparency tools may be misinterpreted and make these uses less possible, if not complemented by other solutions (Case 5).
@@
Alice wants to find and exploit vulnerabilities on a website to demand a ransom payment: She uses a wide range of AI agents to analyze the site’s source code and to test out different sequences of unusual inputs. Because the agents interact primarily with the website through mouse-clicks and simple text-field inputs, synthetic content transparency tools are largely unhelpful in this case.142142142Synthetic content transparency tools will not address mouse-clicks and other forms of browser navigation, which are action-based rather than content-based. A website could attempt to look for watermarks in content submitted to its text fields, but watermarking tends to be unreliable for short pieces of text. Moreover, the site receives only a very small portion of content that the AI has generated; it cannot analyze the AI’s private chain-of-thought in planning and executing this abuse.
@@
Bob lost the ability to speak with much fluency after a medical incident, but he was able to recover his ability to speak by using an AI tool. Now, Bob can once again speak with others over the phone—even with his own voice. Unfortunately, many companies have started to automatically flag and filter out callers they suspect are voiced by AI, given a surge of AI-powered phone scammers. Because Bob’s voice is in fact AI-generated, he is frequently flagged in these settings and has a harder time when he needs to call in for customer support. Without some alternative method, Bob struggles to prove that he is a real person—using AI to help him express himself—rather than a bad actor using AI to deceive.
@@ C.1   Methods to achieve a one-per-person credential limit
How can an issuer authenticate the existing form of ID, so that they do not accidentally issue personhood credentials based on a spoofed document? Some forms of existing ID are already digitally authenticatable today, whereas others may need to be physically presented in person to avoid spoofing.144144144Some national IDs hold near-field communication (NFC) readable, cryptographically signed information like name, address, and date of birth as well as face images. For a description of NFC in ePassports, see [128]. This allows a holder to demonstrably confirm their ownership of the physical card in the digital domain, unlike ID cards in countries like the United States (including some REAL ID-compliant documents), which often rely upon sending a (spoofable) photo or video for digital authentication (depending on the permissions of the verifying party). One challenge may be how to authenticate IDs from different jurisdictions, in a relatively uniform manner. Usefully, there are already global standards for some forms of ID (such as the United Nations’ oversight of ePassport standards), with public key infrastructure built for authenticating these; third-parties can then explore layering zero-knowledge technology atop the documents.145145145For more on ePassports, see [128]. For an overview of how various systems come together to enable identification for cross-border movement, see [73]. For an example of layering zero-knowledge technologies atop these documents, see [287].
@@
Web-of-Trust (WoT) is yet another method for limiting the number of credentials per person: issuers depend on social graph analysis to establish that a person is real (e.g., “Are they vouched for by a person who is strongly believed to be real?”) and has not already received a credential. Often, these systems are “seeded” with initial trusted users from some other method of trust (e.g., existing IDs, biometrics, or in-person “pseudonym parties”) [99]. Someone without a seed-authenticated profile can become authenticated through sufficient vouches from already-trusted parties. While WoT can be a strong method for validating that a person exists, these systems can struggle to confirm uniqueness: A person may be able to get multiple credentials from multiple distinct social circles.152152152Determining whether a person has already received a credential can be a more difficult challenge for WoT-based methods. One way to assess this may be to analyze “Does their social graph look similar to cases where people have been found to have improperly obtained multiple credentials?”, though this will invariably be prone to errors. For a review of WoT-based approaches, see [234]. The most fraud-resistant mechanisms for WoT may involve layering another method, such as biometrics, atop WoT, though in a way such that users still feel comfortable with privacy assurances.153153153For instance, one option for layering simple biometrics atop WoT may include people having a private profile picture, which is viewable only to a small subset of trusted contacts but can still be compared against other private profile pictures in a privacy-preserving way. Use of any biometrics may undermine a system’s reasons for having initially selected WoT as a method, however. To aid in enforcing one-per-person credential limits, WoT issuers may also consider other signals beyond one’s social graph—such as geographic consistency, transaction patterns, and device usage (all commonly used in modern anti-fraud frameworks today).
@@ D.2   Choices for service providers and issuers in an ecosystem with multiple issuers
For a service provider, decisions will often center around which issuers’ PHCs to accept and how these PHC options relate to one another. Trade-offs will often entail balancing the service’s ability to reduce scaled deception by ensuring each person has only one credential, compared with the implications of having fewer choices for its users (such as users feeling pressure to use a credential that is difficult for them to obtain, or the increased concentration of power in a chosen issuer, which creates challenges as discussed in Section 4).